OpenStudy (anonymous):

need help please Find the slope and the y-intercept of the equation y − 3(x − 1) = 0. A.slope = -3 and y-intercept = -3 B.slope = 3 and y-intercept = 3 C.slope = -3 and y-intercept = 3 D.slope = 3 and y-intercept = -3

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So lets try putting the equation into y = mx + b form. Do you know how to do that ?

OpenStudy (texaschic101):

y - 3(x - 1) = 0 --- first we distribute the -3 through the parenthesis y - 3x + 3 = 0 -- now we will subtract 3 from both sides y - 3x = -3 -- now we can add 3x to both sides y = 3x - 3 now, the slope(m) = 3 and the y intercept(b) = -3 any questions ?